OpenStudy (anonymous):

[6.02] Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3)

OpenStudy (jackellyn):

Solve for y in the first equation y= 11-4x Substitute into second equation x +2(11-4x) =8 Simplify/Distribute x + 22 - 8x =8 -7x = -14 x = 2 Put back in first equation to get y 4(2) + y =11 8 +y = 11 y = 3 To check: substitute the values you found in either equation 4(2) +3=11 8+3 = 11 3+2(3)=8 2 + 6 =8

OpenStudy (jackellyn):

So the answer is (x,y)=(2,3)

OpenStudy (anonymous):

thanks

OpenStudy (jackellyn):

No problem :)